Find the probability of the indicated event if ​P(E)equals0.30 and ​P(F)equals0.50.
1
Expert's answer
2020-04-15T14:01:20-0400

Find the probability of the indicated event if P(E) = 0.30 and P(F) = 0.50.

Find P(E or F) if P(E and F) = 0.10


"P(E\ or\ F)=P(E)+P(F)-P(E\ and\ F)"

"P(E\ or\ F)=0.30+0.50-0.10=0.70"